OpenStudy (anonymous):

List the first 3 members of the following sequences defined by (a) tn=2^n+1, (b) tn=3×5^-n please show working out for a and b. Will reward medal :)

OpenStudy (aum):

\[ \large t_n = 2^n + 1 \\ \large t_1 = 2^1 + 1 = 2 + 1 = 3 \\ \large t_2 = ? \\ \large t_3 = ? \]